Find the values of k for each of the following quadratic equations, so that they have two equal roots. (i) 2x^ 2 + kx + 3 = 0

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2x^2+kx+ 3 = 0

On Comparing the given equation withax^2+bx+c= 0

a= 2,b= k andc= 3

As we know, Discriminant =b^2- 4ac

= (k)^2- 4(2) (3)
=k^2- 24
For equal roots, we know, Discriminant = 0
k^2- 24 = 0
k^2= 24
k = ±√24= ±2√6

Thanks